OpenStudy (anonymous):

The amount of carbon-14 present in animal bones t years after the animals death is given by P(t)=P0e^-0.00012097t. how old is an ivory tusk that has lost 17% of it's carbon-14? in years.

OpenStudy (anonymous):

since it has lost 17% , how much percentage is remaining?

OpenStudy (wolf1728):

Half-Life Formulas located here: http://www.1728.org/halflife.htm time = half life * log (bgng amt/endng amt) / log (2) Half life of carbon 14 is 5,730 years Let's say initial amount = 100% and present amount = 83% time = 5,730 * log (100/83) / log (2) time = 5,730 * 0.0809219076 / 0.3010299957 time = 5,730 * 0.2688167584 time = 1,540.32 years